# Is Quadratic Gravity ghost free because the ghost is only virtual?

### v1

@@ Abstract
In [1], inspired by the works [7]-[30], a prescription for calculating the correlation functions in Quadratic Gravity [31]-[32] was presented and further exploited in [2]-[5]. This procedure does not enter in conflict with unitarity. The corresponding Hamiltonian quantization is based on a covariant and contra-variant distinction in the non positive definite metric in the space of states [13]. The Gauss-Ostrogradsky method for higher order theories defines two momentum densities $P_{1}$ and $P_{2}$ and two coordinate densities $Q_{1}$ and $Q_{2}$, one pair is standard, the other ghost like. The approach in [1] involves the continuation $P_{2}\to iP_{2}$ and $Q_{2}\to iQ_{2}$ of the ghost variables acting on kets $|>$ after taking mean values. In the present work, following [42], the LSZ rules are derived, with a formalism adapted to full quartic or higher order theories. The derivation of these rules is possible only after the second quantization of the free theory. The hypothesis for quantization are that $Q_{1},Q_{2}]=\text{"gauge terms"}$, $P_{1},P_{2}]=\text{"gauge terms"}$ and $P_{1},Q_{1}]=iI+\text{"gauge terms"}$. This alone leads to the conclusion that $P_{2},Q_{2}]=-iI+\text{"gauge terms"}$, thus this last pair of variables is ghost like. The graviton contains a massless mode, which is the standard graviton, plus two massive modes with masses $m_{2}$ and $m_{3}$. The third mode is usually interpreted as a ghost in the literature [32]. Here it is shown that, even after making the continuation $P_{2}\to iP_{2}$ and $Q_{2}\to iQ_{2}$, the creation and annihilation operators for this mode commute, thus this particle becomes classical, therefore are not allowed as external lines in a quantum context due to the uncertainty principle. More precisely, the third mode does not appear as a free wave, although it may affect spherically symmetric solutions [32]-[35]. This does not invalidate the model. The effective action $\Gamma$ can be calculated following [31], and can be constrained by the Slavnov-Taylor identities [37]-[38]. The surprising outcome is that the LSZ rules do not contain the third mode as external state, although the correlation functions are conditioned by this mode since it depends on its mass scale. The second mode however, becomes a ghost and is taken care with the methods of [1].
@@ 1. Introduction
Stelle’s gravity [31]-[32] is a renormalizable gravity model, which contains a massless mode and two massive ones. The renormalization program for this model was presented by using the Slavnov-Taylor quantization method [37]-[38]. However, one of the massive modes contributes to the kinetic energy with a wrong sign, and was interpreted as a ghost which spoils unitarity. The theory is of higher order, its equations of motion are of fourth order and its Cauchy problem is well posed [40]-[41]. The unitarity problem was already pointed out as the main issue which obstructs it as a quantum gravity model, at least when the quantization is done around a flat background.
@@
On the other hand, starting from the seminal works of Dirac and Pauli [6]-[7], several authors [8]-[30] and more recently [36] started different programs for quantization of higher order theories and avoiding ghosts. These works are not only related to Quantum Gravity, instead they present different motivations, some of them [8]-[9] are even related to Statistical Physics. All of them bring interesting features which are not evident at first sight. An interesting approach was due to Boulware and Gross [13], who made a formal description of the non positive defined metric in the Hilbert space. These authors consider states such that their inner product $n|m>=\eta_{mn}$ is not positive definite. This reference describes a covariant $\eta_{nm}$, contra-variant $\eta^{nm}$ and mixed metric $\delta_{n}^{m}$. The mixed one is positive definite and is the one to be employed for calculating probabilities. The Hilbert space itself has covariant and contra variant components. The mean values have to be adapted to the covariant or contra variant Hilbert space, but usually negative and non bounded from below quantities become bounded by this covariant and contra-variant distinction. These authors also present several useful formulas for the path integral for these theories, together with some interesting insights. In particular, that depending on the use of the covariant, contra-variant or mixed version, a self adjoint observable may be presented in terms of non hermitian matrices. Despite these interesting findings, those authors open the possibility that the path integral they define may not exist. The main point is the possible presence of real and unbounded exponentials in the path integral, which make the full result divergent.
@@
On the other hand, the authors [4]-[29] made several efforts to physically interpret the results of these references. In particular, in [1] the author made the remarkable observation that the quantization scheme of Boulware-Gross applied to the Stelle model leads, in the euclidean setting, to a perfectly defined path integral. This was further exploited in [2]-[5]. The potentially dangerous exponentials become imaginary and the full result is well defined (in a path integral sense). The main observation is that, in a Hamiltonian setting, there are variables which behave normally under time reversal, and there are ones that not. It is not possible to state in brief the results of these detailed works, but it is worthy to mention that the quantization scheme with these variables, denoted as $P_{2}$ and $Q_{2}$ is consistent after making continuation to imaginary values $P_{2}\to iP_{2}$ and $Q_{2}\to iQ_{2}$. Heuristically, if the commutator $P_{2},Q_{2}]$ has a wrong sign, this continuation removes the unwanted minus sign.
@@
Furthermore, both the results of [13] and [1], strongly point out that this quantization does not enter in any conflict with unitarity. As the final result is the euclidean version of Stelle gravity, one may calculate the physical quantities in the euclidean setting, and then go to the Lorenzian one by continuation to real times. The results will be renormalizable, and is possible that unitarity will be preserved.
@@
The results of all the references mentioned above are of significant interest. However, a full presentation requires an analysis of the LSZ formulas which, in this context, have to be taken with care. The point is that, in this quantization method, the creation and annihilation algebra may have some particularities not present in an standard quantum theory. A program about this topic was initiated in [42]. The LSZ rules require a second quantization of the free model consistent with the principles employed in all these references. In the present work, the idea is to present an LSZ deduction fully adapted to a fourth order theory, regardless how the modes are decomposed. The outcome is very interesting. The usually mode denominated as ghost is a ghost, in the sense that is absent as an external line. This conclusion follows using mild assumptions, which can be relaxed. If they are relaxed a bit, this particle emerges in the physical spectrum, but as a decent state with good properties. In any case, all these features will be discussed along the text.
@@
Furthermore, there is nothing wrong with the covariant and contra-variant formalism of [13]. The presentation here is not a critic against this formalism, it just a way for presenting the results in more concise manner. At the end, the references [1]-[5], [13] and [28]-[29] contain the full formalism, and can be consulted if some clarifications are needed.
The present work is organized as follows. Section 2 presents the arguments showing that the third massive graviton mode in Stelle gravity is either confined or purely virtual at classical and quantum level. The Gupta-Bleuler quantization is taken into account. Section 3 reviews the Gauss-Ostrogradsky quantization for the Pais-Uhlenbeck oscillator, and how the ghost emerges. A simple prescription for avoiding the effect of these ghosts is presented. I believe that this prescription is equivalent to the one presented in [1], and is heavily inspired in that reference and also by [13]. The presentation is intended to be more compact, and this has a price of containing less details. In Section 4 the LSZ rules for the Pais-Uhlenbeck oscillator are derived, in a formalism fully adapted for higher order theories. Section 5 reviews the notable result of reference [1] about the finiteness of the path integral when the above described quantization is presented, which presents a quantization of the model which does not allow divergences due to exponential integrations in the path integral and, at the end, results in the euclidean version of Stelle gravity. In Section 6 the LSZ formulas adapted to this fourth order model is derived, the universal form of the effective action is specified and the effect of the non standard creation annihilation operators is clarified in the last part of this section. Section 7 contains possible research programs related to these results.
2. The classical and quantum disappearance of the third massive graviton mode
@@ 2.1 The free Stelle action in the De Donder gauge
$S_{s}=-\frac{2}{\kappa^{2}}\int\bigg[\frac{1}{2}h^{\mu\nu}\bigg(1+\frac{\kappa^{2}\alpha}{4}\Box\bigg)[\square h_{\mu\nu}-2\partial^{\gamma}\partial_{(\nu}h_{\mu)\gamma}]$
$\frac{1}{2}h\bigg(1+\frac{\kappa^{2}(4\beta-\alpha)}{4}\square\bigg)[2\partial_{\alpha}\partial_{\beta}h^{\alpha\beta}-\square h]-\frac{\kappa^{2}}{2}(\beta-\frac{\alpha}{2})h^{\mu\nu}\partial_{\mu}\partial_{\nu}\partial_{\alpha}\partial_{\beta}h^{\alpha\beta}\bigg]d^{4}x.$ (2.1)
@@
$S_{s}=-\frac{2}{\kappa^{2}}\int\bigg[\frac{1}{2}h^{\mu\nu}\bigg(1+\frac{\kappa^{2}\alpha}{4}\Box\bigg)[\square h_{\mu\nu}-2\partial^{\gamma}\partial_{(\nu}h_{\mu)\gamma}]$
$\frac{1}{2}h\bigg[\frac{2}{3}\bigg(1+\frac{\kappa^{2}(3\beta-\alpha)}{2}\square\bigg)+\frac{1}{3}\bigg(1+\frac{\kappa^{2}\alpha}{4}\square\bigg)\bigg][2\partial_{\alpha}\partial_{\beta}h^{\alpha\beta}-\square h]$
$h^{\mu\nu}\bigg[\bigg(1+\frac{\kappa^{2}\alpha}{4}\partial_{\mu}\partial_{\nu}\bigg)-\bigg(1+\frac{\kappa^{2}(3\beta-\alpha)}{2}\partial_{\mu}\partial_{\nu}\bigg)\bigg]\partial_{\alpha}\partial_{\beta}h^{\alpha\beta}\bigg]d^{4}x.$ (2.2)
@@
$\bigg(1+\frac{\kappa^{2}\alpha}{4}\Box\bigg)\square h_{\mu\nu}-\eta_{\mu\nu}\bigg[\frac{2}{3}\bigg(1+\frac{\kappa^{2}(3\beta-\alpha)}{2}\square\bigg)+\frac{1}{3}\bigg(1+\frac{\kappa^{2}\alpha}{4}\square\bigg)\bigg]\square h$
$2\bigg(1+\frac{\kappa^{2}\alpha}{4}\Box\bigg)\partial^{\gamma}\partial_{(\nu}h_{\mu)\gamma}+\eta_{\mu\nu}\bigg[\frac{2}{3}\bigg(1-\frac{\kappa^{2}(3\beta-\alpha)}{2}\square\bigg)+\frac{1}{3}\bigg(1-\frac{\kappa^{2}\alpha}{4}\square\bigg)\bigg]\partial_{\alpha}\partial_{\beta}h^{\alpha\beta}$
$\bigg[\frac{2}{3}\bigg(1+\frac{\kappa^{2}(3\beta-\alpha)}{2}\square\bigg)+\frac{1}{3}\bigg(1+\frac{\kappa^{2}\alpha}{4}\square\bigg)\bigg]\partial_{\mu}\partial_{\nu}h$
@@
The classical graviton field can be Fourier expanded as in (2.12). Now comes the problem of gauge fixing the model. A customary gauge employed in cosmology is the De Donder gauge
$\partial_{\mu}h^{\mu\nu}=\frac{1}{2}\eta^{\mu\nu}\partial_{\mu}h.$ (2.4)
This gauge will be considered below. But for an easy visualization of results, consider first the Lorenz gauge $\partial_{\mu}h^{\mu\nu}=0$ applied to the equations of motions. The De Donder gauge will be applied later on. The equations in the Lorenz gauge are
$\bigg(1+\frac{\kappa^{2}\alpha}{4}\Box\bigg)\square h_{\mu\nu}-\eta_{\mu\nu}\bigg[\frac{2}{3}\bigg(1+\frac{\kappa^{2}(3\beta-\alpha)}{2}\square\bigg)+\frac{1}{3}\bigg(1+\frac{\kappa^{2}\alpha}{4}\square\bigg)\bigg]\square h$
$\bigg[\frac{2}{3}\bigg(1+\frac{\kappa^{2}(3\beta-\alpha)}{2}\square\bigg)+\frac{1}{3}\bigg(1+\frac{\kappa^{2}\alpha}{4}\square\bigg)\bigg]\partial_{\mu}\partial_{\nu}h=0.$
@@
$\bigg(1+\frac{\kappa^{2}(3\beta-\alpha)}{2}\square\bigg)\square h=0.$
@@
$\bigg(1+\frac{\kappa^{2}\alpha}{4}\Box\bigg)\square h_{\mu\nu}-\frac{1}{3}\eta_{\mu\nu}\bigg(1+\frac{\kappa^{2}\alpha}{4}\square\bigg)\square h$
$\bigg[\frac{2}{3}\bigg(1+\frac{\kappa^{2}(3\beta-\alpha)}{2}\square\bigg)+\frac{1}{3}\bigg(1+\frac{\kappa^{2}\alpha}{4}\square\bigg)\bigg]\partial_{\mu}\partial_{\nu}h=0.$ (2.5)
@@
The task is now to constraint the polarizations in (2.12). For this, note that the Lorenz gauge condition implies that
@@
The second one follows since $h_{1}$ and $h_{3}$ have different Fourier expansions and have to be considered as independent. But this implies that $\epsilon_{1}=0$. Contracting the first with $k^{3\mu}$ and taking into account that $\epsilon^{3}_{\mu\nu}k^{3\nu}=0$ shows that
$\frac{2}{3\omega_{3}(k)}k_{\mu}m_{3}^{2}\bigg(1-\frac{\kappa^{2}\alpha}{4}m_{3}^{2}\bigg)\epsilon_{3}=0,$ (2.7)
thus $\epsilon_{3}=0$ as well. Therefore $h=0$ and the polarization $\epsilon^{i}_{\mu\nu}$ is orthogonal to $k_{i}^{\mu}$ (this is a bit loosely speaking, as the definition of orthogonality in four dimensions is subtle). If the direction of propagation is $\hat{z}$ it is known that the polarizations for the massless mode are
@@
$\Lambda^{\mu}_{\nu}=\frac{1}{m}\left(\begin{array}[]{cccc}\omega&0&0&k\\ 0&1&0&0\\ 0&0&1&0\\ k&0&0&\omega\end{array}\right),$ (2.10)
@@
There is one further important point. All the above discussion has shown that $h=0$, therefore (2.5) becomes
$\bigg(1-\frac{\kappa^{2}\alpha}{4}\Box\bigg)\square h_{\mu\nu}=0.$ (2.11)
But this implies that $h^{3}_{\mu\nu}=0$, since its corresponding mass $m_{3}$ is not present in this equation. Therefore (2.12) becomes
$h^{\mu\nu}(x,t)=\sum_{p}\int\frac{d^{3}k}{(2\pi)^{\frac{3}{2}}\omega_{1}(k)}[\epsilon_{1}^{p\mu\nu}(k)e^{i\omega_{1}(k)t-ik\cdot x}+\epsilon_{1}^{\ast p\mu\nu}(k)e^{-i\omega_{1}(k)t+ik\cdot x}]$
$\sum_{p}\int\frac{d^{3}k}{(2\pi)^{\frac{3}{2}}\omega_{2}(k)}[\epsilon_{2}^{p\mu\nu}(k)e^{i\omega_{2}(k)t-ik\cdot x}+\epsilon_{2}^{\ast p\mu\nu}(k)e^{-i\omega_{2}(k)t+ik\cdot x}].$ (2.12)
This leads to the strange conclusion that the third mode $h^{3}_{\mu\nu}$ is gone as a plane wave.
Assume that the above gauge fixing procedure is the problem for finding this unusual conclusion. Choose instead a healthier way to impose gauge conditions such as the De Donder gauge (2.4). One has to be sure that the above conclusion is not arising due to a bad fixing procedure. The equations of motion (2.3) in this gauge are
$\bigg(1+\frac{\kappa^{2}\alpha}{4}\Box\bigg)\square h_{\mu\nu}-\eta_{\mu\nu}\bigg[\frac{2}{3}\bigg(1+\frac{\kappa^{2}(3\beta-\alpha)}{2}\square\bigg)+\frac{1}{3}\bigg(1+\frac{\kappa^{2}\alpha}{4}\square\bigg)\bigg]\square h$
$2\bigg(1+\frac{\kappa^{2}\alpha}{4}\Box\bigg)\eta_{\gamma(\mu}\partial^{\gamma}\partial_{\nu)}h+\frac{1}{2}\eta_{\mu\nu}\bigg[\frac{2}{3}\bigg(1-\frac{\kappa^{2}(3\beta-\alpha)}{2}\square\bigg)+\frac{1}{3}\bigg(1-\frac{\kappa^{2}\alpha}{4}\square\bigg)\bigg]\square h$
$\bigg[\frac{2}{3}\bigg(1+\frac{\kappa^{2}(3\beta-\alpha)}{2}\square\bigg)+\frac{1}{3}\bigg(1+\frac{\kappa^{2}\alpha}{4}\square\bigg)\bigg]\partial_{\mu}\partial_{\nu}h-\kappa^{2}(\beta-\frac{\alpha}{2})\square\partial_{\mu}\partial_{\nu}h=0.$ (2.13)

### v4

# Scattering amplitudes in Quadratic Gravity in a general formalism

@@ Abstract
In [1], inspired by the works [7]-[30], a prescription for calculating the correlation functions in Quadratic Gravity [31]-[32] was presented and further exploited in [2]-[5]. By construction, it is likely that this procedure does not enter in conflict with unitarity. The corresponding Hamiltonian quantization is based on a covariant and contra-variant distinction in the non positive definite metric in the space of states [13]. The Gauss-Ostrogradsky method for higher order theories defines two momentum densities $P_{1}$ and $P_{2}$ and two coordinate densities $Q_{1}$ and $Q_{2}$, one pair is standard, the other ghost like. The approach in [1] involves the continuation $P_{2}\to iP_{2}$ and $Q_{2}\to iQ_{2}$ of the ghost variables acting on kets $|>$ after taking mean values. In the present work, following [42], the LSZ rules are derived, with a formalism adapted to full quartic or higher order theories. The derivation of these rules is possible only after the second quantization of the free theory. This is considered here, in a more general setting than [42]. It includes a Gupta-Bleuler characterization of the physical states. The effective action $\Gamma$ can be calculated following [31], and can be constrained by the Slavnov-Taylor identities [37]-[38].
@@ 1. Introduction
Stelle’s gravity [31]-[32] is a renormalizable gravity model, which contains a massless mode and two massive ones. In the original reference, the renormalization program for this model was achieved in terms the Slavnov-Taylor quantization method [37]-[38]. However, one of the massive modes contributes to the kinetic energy with a wrong sign, and was interpreted as a ghost which spoils unitarity. The theory is of higher order, its equations of motion are of fourth order and its Cauchy problem is well posed [40]-[41]. The unitarity problem was already pointed out as the main issue which obstructs it as a quantum gravity model, at least when the quantization is done around a flat background.
@@
On the other hand, starting from the seminal works of Dirac and Pauli [6]-[7], several authors [8]-[30] and more recently [36] started different programs for quantization of higher order theories, with the intention of avoiding ghosts. These works are not only related to Quantum Gravity, instead they present different motivations, some of them [8]-[9] are even related to Statistical Physics. All of them bring interesting features which are not evident at first sight. An interesting approach was due to Boulware and Gross [13], who made a formal description of the non positive defined metric in the Hilbert space. These authors consider states such that their inner product $n|m>=\eta_{mn}$ is not positive definite. This reference describes a covariant $\eta_{nm}$, contra-variant $\eta^{nm}$ and mixed metric $\delta_{n}^{m}$. The mixed one is positive definite and is the one to be employed for calculating probabilities. The Hilbert space itself has covariant and contra variant components. The mean values have to be adapted to the covariant or contra variant Hilbert space, but usually negative and non bounded from below quantities become bounded by this covariant and contra-variant distinction. These authors [13] also present several useful formulas for the path integral for these theories, together with some interesting insights. In particular, that depending on the use of the covariant, contra-variant or mixed version, a self adjoint observable may be presented in terms of non hermitian matrices. Despite these interesting findings, those authors open the possibility that the path integral they define may not exist [13]. The main point is the possible presence of real and unbounded exponentials in the path integral, which make the full result divergent.
@@
On the other hand, the authors [4]-[29] made several efforts to physically interpret the results of these references. In particular, in [1] the author made the remarkable observation that the quantization scheme of Boulware-Gross applied to the Stelle model leads, in the euclidean setting, to a perfectly defined path integral. This was further exploited in [2]-[5]. The potentially dangerous exponentials become imaginary and the full result is well defined (in a path integral sense). The main observation is that, in a Hamiltonian setting, there are variables which behave normally under time reversal, and there are ones that not. It is not possible to state in brief the results of these detailed works, but it is worthy to mention that the quantization scheme with these variables, denoted as $P_{2}$ and $Q_{2}$ is consistent after the continuation to imaginary values $P_{2}\to iP_{2}$ and $Q_{2}\to iQ_{2}$. Heuristically, if the commutator $a,a^{\dagger}]$ of an state of the mode has a wrong sign, this continuation removes the unwanted minus sign.
@@
Furthermore, both the results of [13] and [1], strongly point out that this quantization does not enter in any conflict with unitarity. The final result of the whole continuation is the euclidean version of Stelle gravity, one may calculate the physical quantities in the euclidean setting, and then go to the Lorenzian one by continuation to real times. The results will be renormalizable, and is possible that unitarity will be preserved.
@@
The results of all the references mentioned above are of significant interest. However, a full presentation requires an analysis of the LSZ formulas which, in this context, have to be taken with care. The point is that, in this quantization method, the creation and annihilation algebra may have some particularities not present in an standard quantum theory. A program about this topic was initiated in [42]. The LSZ rules require a second quantization of the free model consistent with the principles employed in all these references. In the present work, the idea is to present an LSZ deduction fully adapted to a fourth order theory, regardless how the modes are decomposed. A proper hamiltonian treatement, missing in [42] is included here.
@@
Furthermore, there is nothing wrong with the covariant and contra-variant formalism of [13]. The presentation here is not a critic against this formalism, it just a way for presenting the results in more concise manner. At the end, the references [1]-[5], [13] and [28]-[29] contain the full formalism, and can be consulted if some clarifications are needed since they include a large amount of details.
The present work is organized as follows. Section 2 shows the classical modes of the Stelle model. Also, the Gupta-Bleuler quantization is outlined. Section 3 reviews the Gauss-Ostrogradsky quantization for the Pais-Uhlenbeck oscillator, and how the ghost emerges. A simple prescription for avoiding the effect of these ghosts is presented. I believe that this prescription is equivalent to the one presented in [1], and is heavily inspired in that reference and also by [13]. The presentation is intended to be more compact, and this has a price of containing less details. In Section 4 the LSZ rules for the Pais-Uhlenbeck oscillator are derived, in a formalism fully adapted for higher order theories. Section 5 reviews the notable result of reference [1] about the finiteness of the path integral when the above described quantization is presented, which presents a quantization of the model which does not allow divergences due to exponential integrations in the path integral and, at the end, results in the euclidean version of Stelle gravity. In Section 6 the LSZ formulas adapted to this fourth order model is derived, the universal form of the effective action is specified and the effect of the non standard creation annihilation operators is clarified in the last part of this section. Section 7 contains possible research programs related to these results.
2. Classical and quantum aspects of Quadratic Gravity
@@ 2.1 The free Stelle action in the De Donder gauge
$S_{s}=-\frac{2}{\kappa^{2}}\int\bigg[\frac{1}{2}h^{\mu\nu}\bigg(1-\frac{\kappa^{2}\alpha}{4}\Box\bigg)[\square h_{\mu\nu}-2\partial^{\gamma}\partial_{(\nu}h_{\mu)\gamma}]$
$\frac{1}{2}h\bigg(1-\frac{\kappa^{2}(4\beta-\alpha)}{4}\square\bigg)[2\partial_{\alpha}\partial_{\beta}h^{\alpha\beta}-\square h]-\frac{\kappa^{2}}{2}(\beta-\frac{\alpha}{2})h^{\mu\nu}\partial_{\mu}\partial_{\nu}\partial_{\alpha}\partial_{\beta}h^{\alpha\beta}\bigg]d^{4}x.$ (2.1)
@@
$S_{s}=-\frac{2}{\kappa^{2}}\int\bigg[\frac{1}{2}h^{\mu\nu}\bigg(1-\frac{\kappa^{2}\alpha}{4}\Box\bigg)[\square h_{\mu\nu}-2\partial^{\gamma}\partial_{(\nu}h_{\mu)\gamma}]$
$\frac{1}{2}h\bigg[\frac{2}{3}\bigg(1-\frac{\kappa^{2}(3\beta-\alpha)}{2}\square\bigg)+\frac{1}{3}\bigg(1-\frac{\kappa^{2}\alpha}{4}\square\bigg)\bigg][2\partial_{\alpha}\partial_{\beta}h^{\alpha\beta}-\square h]$
$h^{\mu\nu}\bigg[\bigg(1-\frac{\kappa^{2}\alpha}{4}\partial_{\mu}\partial_{\nu}\bigg)-\bigg(1-\frac{\kappa^{2}(3\beta-\alpha)}{2}\partial_{\mu}\partial_{\nu}\bigg)\bigg]\partial_{\alpha}\partial_{\beta}h^{\alpha\beta}\bigg]d^{4}x.$ (2.2)
@@
$\bigg(1-\frac{\kappa^{2}\alpha}{4}\Box\bigg)\square h_{\mu\nu}-\eta_{\mu\nu}\bigg[\frac{2}{3}\bigg(1-\frac{\kappa^{2}(3\beta-\alpha)}{2}\square\bigg)+\frac{1}{3}\bigg(1-\frac{\kappa^{2}\alpha}{4}\square\bigg)\bigg]\square h$
$2\bigg(1-\frac{\kappa^{2}\alpha}{4}\Box\bigg)\partial^{\gamma}\partial_{(\nu}h_{\mu)\gamma}+\eta_{\mu\nu}\bigg[\frac{2}{3}\bigg(1-\frac{\kappa^{2}(3\beta-\alpha)}{2}\square\bigg)+\frac{1}{3}\bigg(1-\frac{\kappa^{2}\alpha}{4}\square\bigg)\bigg]\partial_{\alpha}\partial_{\beta}h^{\alpha\beta}$
$\bigg[\frac{2}{3}\bigg(1-\frac{\kappa^{2}(3\beta-\alpha)}{2}\square\bigg)+\frac{1}{3}\bigg(1-\frac{\kappa^{2}\alpha}{4}\square\bigg)\bigg]\partial_{\mu}\partial_{\nu}h$
@@
The classical graviton field can be Fourier expanded. Now comes the problem of gauge fixing the model. For an easy visualization of results, consider the Lorenz gauge $\partial_{\mu}h^{\mu\nu}=0$ applied to the equations of motions. The De Donder gauge will be applied later on. The equations in the Lorenz gauge are
$\bigg(1-\frac{\kappa^{2}\alpha}{4}\Box\bigg)\square h_{\mu\nu}-\eta_{\mu\nu}\bigg[\frac{2}{3}\bigg(1-\frac{\kappa^{2}(3\beta-\alpha)}{2}\square\bigg)+\frac{1}{3}\bigg(1-\frac{\kappa^{2}\alpha}{4}\square\bigg)\bigg]\square h$
$\bigg[\frac{2}{3}\bigg(1-\frac{\kappa^{2}(3\beta-\alpha)}{2}\square\bigg)+\frac{1}{3}\bigg(1-\frac{\kappa^{2}\alpha}{4}\square\bigg)\bigg]\partial_{\mu}\partial_{\nu}h=0.$
@@
$\bigg(1-\frac{\kappa^{2}(3\beta-\alpha)}{2}\square\bigg)\square h=0.$ (2.4)
@@
$\bigg(1-\frac{\kappa^{2}\alpha}{4}\Box\bigg)\square h_{\mu\nu}-\frac{1}{3}\eta_{\mu\nu}\bigg(1-\frac{\kappa^{2}\alpha}{4}\square\bigg)\square h$
$\bigg[\frac{2}{3}\bigg(1-\frac{\kappa^{2}(3\beta-\alpha)}{2}\square\bigg)+\frac{1}{3}\bigg(1-\frac{\kappa^{2}\alpha}{4}\square\bigg)\bigg]\partial_{\mu}\partial_{\nu}h=0.$ (2.5)
@@
The task is now to constraints the polarizations. For this, note that the Lorenz gauge condition implies that
@@
The second one follows since $h_{1}$ and $h_{3}$ have different Fourier expansions and have to be considered as independent. But this implies that $\epsilon_{1}=0$. Contracting the first with $k^{3\mu}$ and taking into account that leads to nothing. Therefore $h_{2}=h_{1}=0$ and the polarization $\epsilon^{i}_{\mu\nu}$ is orthogonal to $k_{i}^{\mu}$ (this is a bit loosely speaking, as the definition of orthogonality in four dimensions is subtle). If the direction of propagation is $\hat{z}$ it is known that the polarizations for the massless mode are
@@
$\Lambda^{\mu}_{\nu}=\frac{1}{m}\left(\begin{array}[]{cccc}\omega&0&0&k\\ 0&m&0&0\\ 0&0&m&0\\ k&0&0&\omega\end{array}\right),$ (2.9)
@@
The last thing to analyze is the third mode. After setting $h_{1}=h_{2}=0$ one has that
$\bigg(1-\frac{\kappa^{2}\alpha}{4}m_{3}^{2}\bigg)m_{3}^{2}h_{\mu\nu}-\frac{1}{3}\eta_{\mu\nu}\bigg(1-\frac{\kappa^{2}\alpha}{4}m_{3}^{2}\bigg)m_{3}^{2}h$
$\frac{1}{3}\bigg(1-\frac{\kappa^{2}\alpha}{4}m_{3}^{2}\bigg)\partial_{\mu}\partial_{\nu}h=0.$
The polarization is then
$\epsilon^{3}_{\mu\nu}=\frac{1}{3}\bigg(\eta_{\mu\nu}+\frac{k_{\mu}k_{\nu}}{m_{3}^{2}}\bigg)\epsilon.$ (2.10)
Therefore if the direction of propagation is $k_{\mu}=(\omega(k),0,0,k_{3})$ the polarization matrices reduce to a single one given by
$\epsilon^{+}_{\mu\nu}=\frac{1}{3}\left(\begin{array}[]{cccc}-1+\frac{\omega_{3}^{2}}{m_{3}^{2}}&0&0&\frac{k\omega_{3}}{m_{3}^{2}}\\ 0&1&0&0\\ 0&0&1&0\\ \frac{k\omega_{3}}{m_{3}^{2}}&0&0&1+\frac{k^{2}}{m_{3}^{2}}\end{array}\right).$
The field is decomposed as
$h_{in}^{\mu\nu}=\sum_{p}\int\frac{d^{3}k}{(2\pi)^{\frac{3}{2}}\omega(k)}[\epsilon_{1\mu\nu}e^{i\omega(k)t-ik\cdot x}+\epsilon_{1\mu\nu}^{\ast}e^{-i\omega(k)t+ik\cdot x}]$
$\sum_{p}\int\frac{d^{3}k}{(2\pi)^{\frac{3}{2}}\omega_{2}(k)}[\epsilon_{2\mu\nu}e^{i\omega_{2}(k)t-ik\cdot x}+\epsilon^{\ast}_{2\mu\nu}e^{-i\omega_{2}(k)t+ik\cdot x}].$
$\sum_{p}\int\frac{d^{3}k}{(2\pi)^{\frac{3}{2}}\omega_{3}(k)}[\epsilon_{3\mu\nu}e^{i\omega_{3}(k)t-ik\cdot x}+\epsilon^{\ast}_{3\mu\nu}e^{-i\omega_{3}(k)t+ik\cdot x}].$
The sum is over all allowed polarizations, described above. The next topic is how this gauge conditions are imposed after quantization on the physical states.
2.1.1 The Gupta-Bleuler quantization scheme
Consider the non gauged equations of motion without fixing given in (2.19). As before, the Lorenz gauge $\partial_{\mu}h^{\mu\nu}=0$ will be considered. However, both (2.19) and the gauge has to be valid only when mean values between physical states $|\Psi>$ are taken. The field is expanded as
$h^{\mu\nu}=\sum_{p}\int\frac{d^{3}k}{(2\pi)^{\frac{3}{2}}\omega(k)}[a_{\mu\nu}a_{p}(k)e^{i\omega(k)t-ik\cdot x}+a_{\mu\nu}^{\ast}a_{p}^{c}(k)e^{-i\omega(k)t+ik\cdot x}]$
$\sum_{p}\int\frac{d^{3}k}{(2\pi)^{\frac{3}{2}}\omega_{2}(k)}[b_{\mu\nu}b_{p}(k)e^{i\omega_{2}(k)t-ik\cdot x}+b^{\ast}_{\mu\nu}b_{p}^{c}(k)e^{-i\omega_{2}(k)t+ik\cdot x}].$
$\sum_{p}\int\frac{d^{3}k}{(2\pi)^{\frac{3}{2}}\omega_{3}(k)}[c_{\mu\nu}c_{p}(k)e^{i\omega_{3}(k)t-ik\cdot x}+c^{\ast}_{\mu\nu}c_{p}^{c}(k)e^{-i\omega_{3}(k)t+ik\cdot x}].$
Here $a_{p}(k)$, $b_{p}(k)$ and $c_{p}(k)$ are annihilation operators, and the ones with the upperindex ”c” are the creation ones. The notation ”c” replaces $\dagger$} in order to leave the possibility for a non standard creation-annihiliation algebra.
The sum above is over all the polarizations, not only the allowed by the gauge conditions. The physical states are the ones which will restrict the sums. Therefore the Gupta-Bleuler request is
$\partial^{+}_{\mu}h^{\mu\nu}|\Psi>=0.$
This implies that
Here $\lambda$ included all the polarizations, not only (2.7) and (2.8). The last gives a linear relation between the annihilation operators corresponding to these additional polarizations, whose action destroys the physical states. The averaged equations of motion are now